10x-15y=45
8x-20y=-40

Answer by vasumathi(46)

10x-15y=45
divide this on both sides by 5
2x-3y=9
8x-20y=-40
divide this both sides by 4
2x-5y=-10
2x-3y=9
subtract these two
-2y = -19
y =-19/-2= 9.5
plug in this in the first equation above
that is
2x-3y = 9
2x-3*9.5 = 9
2x-28.5 = 9
2x = 9+28.5= 37.5
x = 37.5/2= 18.75